OpenStudy (anonymous):

A triangle has one angle that measures74 degress. the secound angle is the half the measure of the first angle. what is the measure of the third angle ?

OpenStudy (anonymous):

all the angles equal 180, and you know two angles equal 74 and 37. if we caal the last angle x, then x = 180 - 74 - 37 = 180 - 111 = 69, so the last angle is 69 degrees